Use the method of substitution to solve the system of linear equations.
4a=6-b
4a=2-b
Could you please show me this in detail. I just don't understand it. Thank you

Answer by jim_thompson5910(35256)

4a is both 6-b and 2-b. So 6-b = 2-b

-------------------------------------------------------

Another way to say it is...

4a=6-b ... Start with the first equation.

2-b=6-b ... Replace '4a' with '2-b' (since 4a = 2-b)

-------------------------------------------------------

Now solve for b. Then use the solution to find 'a'.